OpenStudy (kayders1997):

At the limit approaches 0 the limit equals (tanx-x)/x^3

OpenStudy (kayders1997):

@freckles

OpenStudy (freckles):

k we can use l'hospital again if you want

OpenStudy (freckles):

since we have 0/0 when plugin in 0

OpenStudy (kayders1997):

Yes I like that rule

OpenStudy (freckles):

did you know how to diff (tan(x)-x) and x^3?

OpenStudy (kayders1997):

tanx is secx I think?!?! than -x=-1 So (secx-1)/3x^2 I think?

OpenStudy (freckles):

sec^2(x)=(tan(x))'

OpenStudy (freckles):

\[\lim_{x \rightarrow 0}\frac{\sec^2(x)-1}{3x^2}\]

OpenStudy (kayders1997):

Okay is tan the only one that is ^2?

OpenStudy (freckles):

what do you mean?

OpenStudy (kayders1997):

It's been so long I feel like since we did the basic differentiation

OpenStudy (kayders1997):

like tan is sec^2x is that the only one that has the squared in it?

OpenStudy (freckles):

the derivative of cot(x) is -csc^2(x)

OpenStudy (kayders1997):

Okay, ill write that down

OpenStudy (freckles):

derivative of sin(x) is cos(x) derivative of cos(x) is -sin(x) derivative of tan(x) is sec^2(x) derivative of csc(x) is -csc(x) cot(x) derivative of sec(x) is sec(x) tan(x) derivative of cot(x) is -csc^2(x)

OpenStudy (kayders1997):

Omg thank you! :)

OpenStudy (freckles):

but anyways we have 0/0 again when pluggin in 0 so you can do another round of l'hospital

OpenStudy (freckles):

the only thing I see you might have trouble with is diff sec^2(x) you will need chain rule for this

OpenStudy (freckles):

sec^2(x) is the same as writing (sec(x))^2

OpenStudy (freckles):

\[\frac{d}{dx}(\sec(x))^2=2(\sec(x))^{2-1} \cdot (\sec(x))'\]

OpenStudy (kayders1997):

okay so there will be 2(sec(x))*sec(x)tan(x)

OpenStudy (freckles):

that is right or 2 sec^2(x) tan(x)

OpenStudy (freckles):

\[\lim_{x \rightarrow 0} \frac{2 \sec^2(x) \tan(x)}{6x}\]

OpenStudy (freckles):

we could do another round of l'hospital or rearrange this one

OpenStudy (freckles):

your call

OpenStudy (kayders1997):

lets just do another round

OpenStudy (kayders1997):

than it will help me with the basics again

OpenStudy (freckles):

\[\lim_{x \rightarrow 0} \frac{2 \sec^2(x) \frac{\sin(x)}{\cos(x)}}{6x} \\ =\lim_{x \rightarrow 0} \frac{2 \sec^2(x)}{\cos(x)} \frac{\sin(x)}{6x} \\ =\frac{2}{6}\lim_{x \rightarrow 0} \frac{ \sec^2(x)}{\cos(x)} \frac{\sin(x)}{x}\]

OpenStudy (freckles):

see if you can manage the limit from here

OpenStudy (kayders1997):

wait it you did DS wouldn't your answer just be 2/6 or 1/3?

OpenStudy (kayders1997):

*if

OpenStudy (freckles):

2/6 reduces to 1/3 yes you have just after my last line the following: \[\frac{2}{6} \cdot \frac{1^2}{1} \cdot 1 =\frac{2}{6}=\frac{1}{3}\]

OpenStudy (freckles):

as x approaches 0: sin(x)/x goes to 1 sec^2(x) goes to 1 since sec^2(0) is 1 cos(x) goes to 1 since cos(1) is 1 so sec^2(x)/cos(x)=1

OpenStudy (freckles):

oops mistype

OpenStudy (freckles):

cos(0) is 1*

OpenStudy (kayders1997):

Note to jot down lol

OpenStudy (kayders1997):

Thank you so much :)